OpenStudy (anonymous):

A particle in free-fall: takes 42 seconds to travel 490 meters at g= 9.8 m/s2.

OpenStudy (shamim):

time t=42sec height h=490 m gravitational acceleration g=9.8m/sec^2

OpenStudy (shamim):

now wt do u wanna know

OpenStudy (shamim):

we know h=ut+(1/2)gt^2 490=u*42+(1/2)*9.8*42^2 now u can find out the value of u here u is initial velocity